solve by the elimination method 7x+5y=7 4x-5y=-51
Answer by mananth(16946)

7x+5y=7 .............1
4x-5y=-51.............2

Add the two equations

11x+=-44

/11

x=-4

plug value of x In (1)

7x+5y=7
-28+5y= 7
5y=7+28
5y=35
y=7